Beta Ltd has a product that sells for $50 and is produced at a variable production cost of $24 per unit and a variable non-production cost of $8. The variable production costs can be reduced 25% by installing a new piece of equipment. The Installation of the new equipment will increase fixed costs from the present level of $122,400 to $159,000.Calculate the margin of safety (in units, RM and nearest %) based on the present break-even point and the new break-even point if the budgeted sales are 8,500 units.                


1
Expert's answer
2021-08-09T15:53:11-0400
"TC=FC+VC"

"VC=(24+8)Q"

"50Q=122,400+32Q"

"Q=6,800"

"50Q=159,000+(18+8)Q"

"Q=6,625"